Robie is 10 years older than Micah. In three years, the product of her age and Micah's age five years ago will be 80 years. Find their present ages.
Answer by stanbon(75887)

Robie is 10 years older than Micah. In three years, the product of her age and Micah's age five years ago will be 80 years. Find their present ages.
----
Equations:
r = m + 10
(r+3)(m-5) = 80
----------------------
substitute for "r" and solve for "m":
(m+10+3)(m-5) = 80
(m+13)(m-5) = 80
Solve for m; then for r.
===
Cheers,
Stan H.